A satellite is launched into a circular of radius R around the earth.While a second satellite launched in an orbit of radius 1.01R.The period of the second satellite is longer than the first one by approximately(please express in %)

Grade:12

1 Answers

Chetan Mandayam Nayakar
13 years ago

 

Dear Sunny, 

according to Kepler's third law T2 isproportional to R3. Thus t1=kR3/2,t2=kR3/2(1+0.01)3/2=kR3/2(1+(0.01*1.5)). Thus the time period of the second satellite is longer than that of the first one by approximately 1.5%.

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar respectively : Click here to download the toolbar.. 

Askiitians Expert

Chetan Mandayam Nayakar – IIT Madras